OpenStudy (anonymous):

What is the circumference of this circle? Use π ≈ 3.14. Round your answer to the nearest tenth of a centimeter. A. 69.1 cm B. 22.0 cm C. 34.5 cm D. 17.3 cm

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

here we can apply this formula: \(C= \pi \cdot D= 3.14 \times 11=...?\)

OpenStudy (waterdragon):

he is reported